OpenStudy (across):

So I just heard my advisor lecture his students that performing a double integral on an expression is not the same thing as integrating that expression twice (all with respect to a single, independent variable). The context of the problem was the solution of an autonomous, 2nd order ODE: \[\frac{d^2y}{dx^2}=f(x).\]Apparently, \[\int \int \frac{d^2y}{dx^2}=\int \int f(x)\] is not the same thing as \[\int [ \int \frac{d^2y}{dx^2} ]=\int [ \int f(x) ]\]Why did he say that?

myininaya (myininaya):

i don't know i always work inside to outside i don't see how it isn't the same (yuck double negative) i see it as the same

OpenStudy (across):

He further added, although rather ambiguously, that a second integral had more to do with spatial concepts. I didn't want to question him for fear that my students would lose trust in my reviews during TA sessions, had I been wrong in some remote, arcane sense. Perhaps they're not the same thing when viewed from a more convoluted position?

myininaya (myininaya):

so he didn't have any limits?

myininaya (myininaya):

i remember you can't switch the outside integral with the inside integral sometimes

OpenStudy (across):

There were no limits; the problem was stated just as I wrote it.

myininaya (myininaya):

imran any thoughts?

OpenStudy (anonymous):

no , my brain is melting from physics

myininaya (myininaya):

physics>math ?

OpenStudy (across):

If you know quantum physics, talk to me. xd

OpenStudy (anonymous):

no, just classical

myininaya (myininaya):

ok i'm going to see if i can find anything on your double integral question

OpenStudy (anonymous):

I think we can switch order if limits are constant, or something

OpenStudy (across):

I totally agree with you that they're not the same thing when limits are incorporated. Perhaps he just said that to discourage students from thinking about it that way?

myininaya (myininaya):

f(x)=5x \[\int\limits_{}^{}\int\limits_{}^{}(5x) dx dx=\int\limits_{}^{}(\frac{5x^2}{2}+C)dx=\frac{5x^3}{2(3)}+Cx+D=\frac{5x^3}{6}+Cx+D\] is this incorrect?

OpenStudy (across):

It's correct.

myininaya (myininaya):

but i did this \[\int\limits_{}^{} [\int\limits_{}^{}5x dx] dx\]

myininaya (myininaya):

i performed the inside integral first then the outside

OpenStudy (across):

You know what, I think he may have just messed up because, as far as I know, that's how you solve a double integral.

OpenStudy (zarkon):

what was he integrating with respect to...did he leave that off like you wrote?

OpenStudy (across):

He stated we were talking about one dependent and one independent variable. So the integral (in this case) was (implicitly) with respect to x.

OpenStudy (across):

I'll ask him about this and see what reason he gives me.

OpenStudy (zarkon):

I only see a difference when there are limits dxdy vs dA

myininaya (myininaya):

\[\int\limits_{}^{}\int\limits_{}^{}(5x)dxdy=\int\limits_{}^{}(\frac{5x^2}{2}+C)dy=\frac{5x^2y}{2}+Cy+D\] \[\neq\] \[\int\limits_{}^{}\int\limits_{}^{}(5x)dydx=\int\limits_{}^{}(5xy+C) dx=\frac{5x^2y}{2}+Cx+D\] maybe he left the limits off because he was trying to be general

myininaya (myininaya):

not the limits

myininaya (myininaya):

but what we are integrating with respect too

myininaya (myininaya):

ok i have to figure out how to program a line in c++

myininaya (myininaya):

eventhough imran wants me to use cg

OpenStudy (zarkon):

we need to be a little careful when we do things like... \["\int\limits_{}^{}\int\limits_{}^{}(5x)dxdy=\int\limits_{}^{}(\frac{5x^2}{2}+C)dy=\frac{5x^2y}{2}+Cy+D"\] if we have a two variable function then \[\int\int 5x \,\,dxdy=\int\left(\frac{5}{2}x^2+g(y)\right)dy\] since \[\frac{\partial}{\partial x}\left(\frac{5}{2}x^2+g(y)\right)=5x^2\]

myininaya (myininaya):

you are right

myininaya (myininaya):

gj zarkon! i would give you another medal if i could

OpenStudy (anonymous):

As far as I know, he was wrong. (I don't see any reason the brackets modify the scenario.) I also know that this input is not exactly relevant 10 months later.

OpenStudy (zarkon):

nobody noticed my typo ;) \[\frac{\partial}{\partial x}\left(\frac{5}{2}x^2+g(y)\right)=5x\]